# On the Genericity of the Spectrum Intervalization for Multi-Frequency Quasiperiodic Schrödinger Operators

### v1

@@ Abstract
This paper proves a genericity conjecture by Goldstein, Schlag, and Voda[Invent. Math.217(2019)] for multi-frequency quasiperiodic Schrödinger operators. Specifically, we show that for almost all coefficients of real trigonometric polynomial potentials, the spectrum forms a single interval under strong coupling conditions. This confirms a long-standing intuition by Chulaevky and Sinai[Comm.Math.Phys.125(1989)] that the spectrum typically intervals for generic potentials, and extends the existence results of Goldstein et al. to a full measure setting. Our proof relies on tools from differential topology, measure theory, and analytic function theory.
@@
Quasiperiodic Schrödinger operators , Spectrum , Genericity , Cartan estimates , Transversality theory
@@ 1 Introduction
In this paper, we prove the genericity conjecture by Goldstein et al. [11], showing that for almost all coefficients of real trigonometric polynomials, the potential belongs to class $\mathfrak{G}$, which ensures the spectrum is a single interval. Our approach combines tools from differential topology, measure theory, and analytic function theory, specifically leveraging the parametric transversality theorem, and Cartan-type estimates to establish the full-measure property of class $\mathfrak{G}$. This provides a comprehensive framework for understanding the generic behavior of these operators and confirms the original intuition of Chulaevsky and Sinai [8].
@@ 2 Preliminaries
2.1 Class $\mathfrak{G}$ of Potetials
@@
with coefficients $\mathbf{c}=(c_{m})\in\mathbb{R}^{N}$, where $N$ is the number of integer vectors $m\in\mathbb{Z}^{d}$ satisfying the condition $|m|=\sum_{j=1}^{d}|m_{j}|\leq n$. By a combinatorial count, the dimension $N$ is given explicitly by:
@@
This counts the number of solutions to the inequality $|m|\leq n$ in integers.
@@
    $g_{V,h,i,j}(x)=\operatorname{det}\left[\begin{array}[]{cc}\partial_{x_{i}}V(x)&\partial_{x_{j}}V(x)\\ \partial_{x_{i}}V(x+h)&\partial_{x_{j}}V(x+h)\end{array}\right].$
    (see [7] for the origin of such estimates).
@@
Here, $\mathfrak{c}_{1}>0$ is a constant, and the Cartan estimates are based on the theory of analytic functions as developed by Cartan [7] and applied in [11].
@@ 2.2 Statement of Main Theorem
We now state the main theorem, which proves the genericity conjecture by Goldstein et al. [11, Remark 1.2(b)].
@@
The proof of Theorem 2.2 is based on the parametric transversality theorem from differential topology.
@@ 2.3 Parametric Transversality Theorem
For a smooth manifold $X$ and a point $x\in X$, the tangent space $T_{x}X$ is the vector space of all possible directions one can move from $x$ while remaining on the manifold. Intuitively, if $X$ is a surface in $\mathbb{R}^{3}$, then $T_{x}X$ is the plane tangent to $X$ at $x$. The dimension of $T_{x}X$ equals the dimension of $X$ at $x$.
@@
where $df_{x}:T_{x}X\to T_{f(x)}Y$ is the derivative of $f$ at $x$. Transversality means that the image of $df_{x}$ and the tangent space of $Z$ together span the entire tangent space of $Y$, ensuring non-degenerate intersections.
@@
This is a standard result in differential topology. One proof applies Sard’s theorem to the restriction of the projection $\pi:P\times X\to P$ to the submanifold $F^{-1}(Z)$, which is smooth due to the transversality of $F$. The set of parameters $p$ for which $F_{p}$ is not transverse to $Z$ is contained in the set of critical values of this projection, hence has measure zero. For complete details, see [2, Theorem 8.4] or [13, Chapter 3, Theorem 2.7]. ∎
@@
In our context, $X=\mathbb{T}^{d}$ is the torus, $Y=\mathbb{R}^{m}$ for appropriate $m$ (e.g., $m=d+1$ for critical point analysis), and $P=\mathbb{R}^{N}$ is the space of trigonometric polynomial coefficients. The submanifold $Z$ typically represents a set where degeneracies occur (e.g., where critical points become non-Morse). The transversality of $F$ ensures that for generic $\mathbf{c}$, the potential $V(\cdot;\mathbf{c})$ avoids these degeneracies. This is used extensively in Section 3 to prove genericity of Morse properties and unique extrema.
@@ 3 Proof of the Main Results
The proof relies on two key lemmas.
@@
We prove the lemma using the parametric transversality theory, which are well-suited for analyzing generic properties of families of functions. Let
$V(x;\mathbf{c})=\sum_{m\in\mathbb{Z}^{d}:|m|\leq n}c_{m}e^{2\pi im\cdot x}$ (3.1)
be the trigonometric polynomial potential, where $x\in\mathbb{T}^{d}$ and $\mathbf{c}\in\mathbb{R}^{N}$. We treat $V$ as a smooth (in fact, analytic) function in both $x$ and $\mathbf{c}$.
@@
A function $V(\cdot;\mathbf{c})$ is Morse if all its critical points are non-degenerate, i.e., for every $x$ such that $\nabla_{x}V(x;\mathbf{c})=0$, the Hessian matrix $H(V)(x;\mathbf{c})$ is non-singular. To prove that the set
$\mathcal{B}^{(i)}=\{\mathbf{c}\in\mathbb{R}^{N}\mid V(\cdot;\mathbf{c})\text{ is not Morse}\}$
has measure zero, we employ the Parametric Transversality Theorem. Consider the map
$\Phi:\mathbb{T}^{d}\times\mathbb{R}^{N}\to\mathbb{R}^{d}\times\mathbb{R},\quad\Phi(x,\mathbf{c})=\left(\nabla_{x}V(x;\mathbf{c}),\det(H(V)(x;\mathbf{c}))\right).$
A point $(x,\mathbf{c})$ is a degenerate critical point if $\Phi(x,\mathbf{c})=(0,0)$. The set of degenerate critical points is $\mathcal{D}=\Phi^{-1}(0,0)$. By Theorem 2.5, if $\Phi$ is transverse to the submanifold $\{0\}\times\{0\}\subset\mathbb{R}^{d}\times\mathbb{R}$ (i.e., the derivative $D\Phi$ is surjective at points in $\mathcal{D}$), then for almost every $\mathbf{c}\in\mathbb{R}^{N}$, the function $V(\cdot;\mathbf{c})$ has no degenerate critical points, meaning it is Morse.
Since $V$ is analytic in $x$ and linear in $\mathbf{c}$, the map $\Phi$ is smooth. To verify the transversality condition, we explicitly compute the derivative $D\Phi$ with respect to $\mathbf{c}$. For fixed $(x,\mathbf{c})$, the derivative $D_{\mathbf{c}}\Phi(x,\mathbf{c}):\mathbb{R}^{N}\to\mathbb{R}^{d}\times\mathbb{R}$ acts on a vector $v=(v_{m})\in\mathbb{R}^{N}$ as follows:
$D_{\mathbf{c}}\Phi(x,\mathbf{c})[v]$ $=\left(D_{\mathbf{c}}(\nabla_{x}V)(x;\mathbf{c})[v],\ D_{\mathbf{c}}(\det H(V))(x;\mathbf{c})[v]\right)$
$=\left(\nabla_{x}\left(\frac{\partial V}{\partial\mathbf{c}}[v]\right),\ \operatorname{tr}\left(\operatorname{adj}(H(V))\cdot D_{\mathbf{c}}H(V)[v]\right)\right),$ (3.2)
where $\frac{\partial V}{\partial\mathbf{c}}[v]=\sum_{m}v_{m}e^{2\pi im\cdot x}$ is the directional derivative of $V$ with respect to $\mathbf{c}$ in the direction $v$, and we used the identity for the derivative of the determinant.
The key observation is that the components of $\nabla_{x}V$ and $\det H(V)$ are linear in $\mathbf{c}$ with coefficients given by the Fourier basis functions and their derivatives. Specifically, for each coordinate $i=1,\dots,d$:
$\frac{\partial V}{\partial x_{i}}=\sum_{m}c_{m}(2\pi im_{i})e^{2\pi im\cdot x},$
and the Hessian determinant $\det H(V)$ is a homogeneous polynomial of degree $d$ in the second derivatives $\frac{\partial^{2}V}{\partial x_{i}\partial x_{j}}$, which are themselves linear in $\mathbf{c}$ with coefficients involving monomials in $m$.
The linear independence of the Fourier basis $\{e^{2\pi im\cdot x}\}$ over $\mathbb{C}$ (and hence over $\mathbb{R}$ for real parts) implies that the set of functions
$\left\{\frac{\partial V}{\partial x_{i}}(\cdot;\mathbf{c}),\det H(V)(\cdot;\mathbf{c})\right\}$
has full rank for generic $x$ when considered as linear functionals in $\mathbf{c}$. More precisely, the variational derivative of $\Phi$ with respect to $\mathbf{c}$ at a point $(x,\mathbf{c})$ is given by the Jacobian matrix whose columns are the partial derivatives with respect to each $c_{m}$. For the first component:
$\frac{\partial}{\partial c_{m}}\left(\frac{\partial V}{\partial x_{i}}\right)=2\pi im_{i}e^{2\pi im\cdot x},$
and for the second component, by the chain rule:
$\frac{\partial}{\partial c_{m}}(\det H(V))=\operatorname{tr}\left(\operatorname{adj}(H(V))\cdot\frac{\partial H(V)}{\partial c_{m}}\right),$
where $\frac{\partial H(V)}{\partial c_{m}}$ is the Hessian of the basis function $e^{2\pi im\cdot x}$. Since the Fourier basis is linearly independent, the vectors
$\left(2\pi im_{i}e^{2\pi im\cdot x}\right)_{i=1}^{d}\quad\text{and}\quad\operatorname{tr}\left(\operatorname{adj}(H(V))\cdot\frac{\partial H(V)}{\partial c_{m}}\right)$
collectively span $\mathbb{R}^{d}\times\mathbb{R}$ as $m$ varies over the index set, for generic $x$ where $H(V)$ is non-degenerate. This follows because the number of independent Fourier modes exceeds the dimension $d+1$ of the target space.
Thus, $D_{\mathbf{c}}\Phi$ is surjective at points where $\Phi(x,\mathbf{c})=(0,0)$, ensuring transversality. By the parametric transversality theorem, the set of $\mathbf{c}$ for which $V(\cdot;\mathbf{c})$ fails to be Morse has measure zero. Hence, $\mathcal{B}^{(i)}$ has measure zero.
@@
We now show that the set
$\mathcal{B}^{(ii)}=\{\mathbf{c}\in\mathbb{R}^{N}\mid V(\cdot;\mathbf{c})\text{ lacks unique global extrema}\}$
has measure zero. A function $V(\cdot;\mathbf{c})$ has unique global extrema if the global maximum and minimum values are attained at unique points in $\mathbb{T}^{d}$. Since $\mathbb{T}^{d}$ is compact, global extrema exist and are critical points. If $V(\cdot;\mathbf{c})$ is Morse (which holds for almost every $\mathbf{c}$ by Part 1), then all critical points are isolated. However, uniqueness of global extrema requires that no two distinct critical points share the same extreme value.
To prove this, we employ a transversality argument on the critical set. Let $\mathcal{C}\subset\mathbb{T}^{d}\times\mathbb{R}^{N}$ be the set of critical points, defined by:
@@
Since $V$ is analytic in $x$ and linear in $\mathbf{c}$, the implicit function theorem implies that $\mathcal{C}$ is a smooth manifold of dimension $N$ (the same as the parameter space) for generic $\mathbf{c}$.
Define the evaluation map on the critical set:
$\Psi:\mathcal{C}\to\mathbb{R},\quad\Psi(x,\mathbf{c})=V(x;\mathbf{c}).$
The global extrema are unique if the restriction of $\Psi$ to each fiber $\mathcal{C}_{\mathbf{c}}=\{x\mid(x,\mathbf{c})\in\mathcal{C}\}$ is injective on the set of extreme values. This is equivalent to requiring that for any two distinct critical points $x\neq y$ in $\mathcal{C}_{\mathbf{c}}$, we have $V(x;\mathbf{c})\neq V(y;\mathbf{c})$.
To analyze this condition, consider the double critical set and the difference map:
@@
which is a smooth manifold of dimension 2Nd2N-d (since the constraint $x\neq y$ removes a submanifold of dimension N1N-1, and $\dim\mathcal{C}=N$). Define the difference map:
@@
The set of parameters $\mathbf{c}$ for which $V(\cdot;\mathbf{c})$ has non-unique extrema is exactly the projection onto $\mathbb{R}^{N}$ of the set:
$\mathcal{E}=\Xi^{-1}(0)=\{((x,\mathbf{c}),(y,\mathbf{c}))\in\mathcal{C}^{(2)}\mid V(x;\mathbf{c})=V(y;\mathbf{c})\}.$
We now show that $\Xi$ is transverse to $\{0\}\subset\mathbb{R}$ (i.e., $D\Xi$ is surjective at points where $\Xi=0$). For a point $((x,\mathbf{c}),(y,\mathbf{c}))\in\mathcal{E}$, the derivative $D\Xi$ acts on tangent vectors $(\dot{x},\dot{\mathbf{c}}_{1},\dot{y},\dot{\mathbf{c}}_{2})\in T_{(x,\mathbf{c})}\mathcal{C}\times T_{(y,\mathbf{c})}\mathcal{C}$ as:
$D\Xi[(\dot{x},\dot{\mathbf{c}}_{1},\dot{y},\dot{\mathbf{c}}_{2})]=\langle\nabla_{x}V(x;\mathbf{c}),\dot{x}\rangle+\langle D_{\mathbf{c}}V(x;\mathbf{c}),\dot{\mathbf{c}}_{1}\rangle-\langle\nabla_{y}V(y;\mathbf{c}),\dot{y}\rangle-\langle D_{\mathbf{c}}V(y;\mathbf{c}),\dot{\mathbf{c}}_{2}\rangle.$
@@
Since $(x,\mathbf{c}),(y,\mathbf{c})\in\mathcal{C}$, we have $\nabla_{x}V(x;\mathbf{c})=0$ and $\nabla_{y}V(y;\mathbf{c})=0$, so the terms involving $\dot{x}$ and $\dot{y}$ vanish. Thus:
$D\Xi[(\dot{x},\dot{\mathbf{c}}_{1},\dot{y},\dot{\mathbf{c}}_{2})]=\langle D_{\mathbf{c}}V(x;\mathbf{c})-D_{\mathbf{c}}V(y;\mathbf{c}),\dot{\mathbf{c}}\rangle,$
where $\dot{\mathbf{c}}=\dot{\mathbf{c}}_{1}=\dot{\mathbf{c}}_{2}$ since $\mathbf{c}$ is the same for both points. The linear independence of the Fourier basis $\{e^{2\pi im\cdot x}\}$ ensures that for generic $x\neq y$, the vectors $D_{\mathbf{c}}V(x;\mathbf{c})$ and $D_{\mathbf{c}}V(y;\mathbf{c})$ are linearly independent. Specifically:
$D_{\mathbf{c}}V(x;\mathbf{c})=\left(e^{2\pi im\cdot x}\right)_{|m|\leq n},\quad D_{\mathbf{c}}V(y;\mathbf{c})=\left(e^{2\pi im\cdot y}\right)_{|m|\leq n},$
and since $x\neq y$, the exponential functions are linearly independent for distinct frequencies. Thus, $D_{\mathbf{c}}V(x;\mathbf{c})-D_{\mathbf{c}}V(y;\mathbf{c})\neq 0$ for generic $x\neq y$, ensuring that $D\Xi$ is surjective onto $\mathbb{R}$.
@@
By Theorem 2.5, the set $\mathcal{E}$ is a smooth submanifold of $\mathcal{C}^{(2)}$ with codimension 1. The projection $\pi:\mathcal{C}^{(2)}\to\mathbb{R}^{N}$ given by $\pi((x,\mathbf{c}),(y,\mathbf{c}))=\mathbf{c}$ is a smooth map. Since $\dim\mathcal{E}=\dim\mathcal{C}^{(2)}-1=2N-d-1$ and $\dim\mathbb{R}^{N}=N$, if $N>d+1$ (which holds for sufficiently high-degree trigonometric polynomials), then $\dim\mathcal{E}<N$, and by Sard’s theorem, the set of critical values of $\pi|_{\mathcal{E}}$ has measure zero in $\mathbb{R}^{N}$. The set of $\mathbf{c}$ for which $V(\cdot;\mathbf{c})$ has non-unique extrema is contained in this set of critical values, hence has measure zero. Therefore, $\mathcal{B}^{(ii)}$ has measure zero.
Conclusion.

### v2

@@ Abstract
This paper proves a genericity conjecture by Goldstein, Schlag, and Voda [Invent. Math. 217 (2019)] for multi-frequency quasiperiodic Schrödinger operators. Specifically, we show that for almost all coefficients of real trigonometric polynomial potentials, the spectrum forms a single interval under strong coupling conditions. This confirms a long-standing intuition by Chulaevsky and Sinai [Comm. Math. Phys. 125 (1989)] that the spectrum typically consists of an interval for generic potentials, and extends the existence results of Goldstein et al. to a full measure setting. Our proof relies on tools from differential topology, measure theory, and analytic function theory.
@@
Quasiperiodic Schrödinger operators , Spectrum , Genericity , Cartan estimates , Transversality theory.
@@ 1 Introduction
In this paper, we prove the genericity conjecture by Goldstein et al. [13], showing that for almost all coefficients of real trigonometric polynomials, the potential belongs to class $\mathfrak{G}$, which ensures the spectrum is a single interval. Our approach combines tools from differential topology, measure theory, and analytic function theory, specifically leveraging the parametric transversality theorem and Cartan-type estimates to establish the full-measure property of class $\mathfrak{G}$. This provides a comprehensive framework for understanding the generic behavior of these operators and confirms the original intuition of Chulaevsky and Sinai [9].
@@ 2 Preliminaries
2.1 Class $\mathfrak{G}$ of Potentials
@@
with coefficients $\mathbf{c}=(c_{m})\in\mathbb{R}^{N}$, where NN is the number of integer vectors $m\in\mathbb{Z}^{d}$ satisfying $|m|=\sum_{j=1}^{d}|m_{j}|\leq n$. A combinatorial count gives
@@
    $g_{V,h,i,j}(x)=\operatorname{det}\begin{bmatrix}\partial_{x_{i}}V(x)&\partial_{x_{j}}V(x)\\ \partial_{x_{i}}V(x+h)&\partial_{x_{j}}V(x+h)\end{bmatrix}$.
    (see [8] for the origin of such estimates). The constant $\mathfrak{c}_{1}>0$ is the same as in [13].
@@ 2.2 Statement of Main Theorem
We now state the main theorem, which proves the genericity conjecture by Goldstein et al. [13, Remark 1.2(b)].
@@
The proof of Theorem 2.2 relies on two key lemmas: Lemma 3.1, which uses the parametric transversality theorem to establish genericity of Morse functions and unique extrema, and Lemma 3.6, which uses Cartan estimates and Borel-Cantelli arguments to show that the required analytic estimates hold for almost every coefficient.
@@ 2.3 Parametric Transversality Theorem
For a smooth manifold $X$ and a point $x\in X$, the tangent space $T_{x}X$ is the vector space of all possible directions one can move from $x$ while remaining on the manifold. The dimension of $T_{x}X$ equals the dimension of $X$ at $x$.
@@
where $df_{x}:T_{x}X\to T_{f(x)}Y$ is the derivative of $f$ at $x$.
@@
This is a standard result in differential topology. One proof applies Sard’s theorem to the restriction of the projection $\pi:P\times X\to P$ to the submanifold $F^{-1}(Z)$, which is smooth due to the transversality of $F$. The set of parameters $p$ for which $F_{p}$ is not transverse to $Z$ is contained in the set of critical values of this projection, hence has measure zero. For complete details, see [2, Theorem 8.4] or [15, Chapter 3, Theorem 2.7]. ∎
@@
In our context, $X=\mathbb{T}^{d}$ is the torus, $Y=\mathbb{R}^{m}$ for appropriate $m$ (e.g., $m=d+1$ for critical point analysis), and $P=\mathbb{R}^{N}$ is the space of trigonometric polynomial coefficients. The submanifold $Z$ typically represents a set where degeneracies occur (e.g., where critical points become non-Morse). The transversality of $F$ ensures that for generic $\mathbf{c}$, the potential $V(\cdot;\mathbf{c})$ avoids these degeneracies.
2.4 Borel-Cantelli Lemma
Proposition 2.7 (Borel-Cantelli Lemma).
Let $(\Omega,\mathcal{F},\mu)$ be a measure space with $\mu(\Omega)<\infty$ (or more generally a probability space).
  1. (1)
    First Borel-Cantelli Lemma: If $\{A_{k}\}_{k\geq 1}$ is a sequence of measurable sets such that $\sum_{k=1}^{\infty}\mu(A_{k})<\infty$, then
    $\mu\Bigl(\bigcap_{n=1}^{\infty}\bigcup_{k=n}^{\infty}A_{k}\Bigr)=0.$
    In words, almost every point belongs to only finitely many $A_{k}$.
  2. (2)
    Second Borel-Cantelli Lemma: If $\{A_{k}\}_{k\geq 1}$ are independent events in a probability space and $\sum_{k=1}^{\infty}\mu(A_{k})=\infty$, then
    $\mu\Bigl(\bigcap_{n=1}^{\infty}\bigcup_{k=n}^{\infty}A_{k}\Bigr)=1,$
    i.e., almost every point belongs to infinitely many $A_{k}$.
For a proof see e.g. [11, Theorem 2.3.1 and 2.3.7] or [3, Theorem 4.2 and 4.3].
@@ 3 Proof of the Main Results
3.1 Genericity of Morse Properties and Unique Extrema
@@
We prove the lemma using algebraic and differential topological methods. Let
$V(x;\mathbf{c})=\sum_{m\in\mathbb{Z}^{d}:|m|\leq n}c_{m}e^{2\pi im\cdot x}$
be the trigonometric polynomial potential, where $x\in\mathbb{T}^{d}$ and $\mathbf{c}\in\mathbb{R}^{N}$. $V$ is smooth (in fact, analytic) in both $x$ and $\mathbf{c}$.
@@
A function $V(\cdot;\mathbf{c})$ is Morse iff it has no degenerate critical points, i.e., there is no pair $(x,\mathbf{c})$ with $\nabla_{x}V(x;\mathbf{c})=0$ and $\det H(V)(x;\mathbf{c})=0$. Define
$\Phi:\mathbb{T}^{d}\times\mathbb{R}^{N}\to\mathbb{R}^{d}\times\mathbb{R},\qquad\Phi(x,\mathbf{c})=\bigl(\nabla_{x}V(x;\mathbf{c}),\;\det H(V)(x;\mathbf{c})\bigr).$
The set of degenerate critical points is $\mathcal{D}=\Phi^{-1}(0,0)$. The projection $\pi(x,\mathbf{c})=\mathbf{c}$ sends $\mathcal{D}$ onto $\mathcal{B}^{(i)}$, the set of parameters for which $V(\cdot;\mathbf{c})$ has at least one degenerate critical point. Thus it suffices to prove that $\pi(\mathcal{D})$ has Lebesgue measure zero.
We use an algebraic argument. Because $V$ is a trigonometric polynomial, the equations $\nabla_{x}V(x;\mathbf{c})=0$ are linear in $\mathbf{c}$ with coefficients that are analytic functions of $x$ (specifically, Fourier exponentials). The equation $\det H(V)(x;\mathbf{c})=0$ is a polynomial of degree $d$ in the second derivatives, hence also polynomial in $\mathbf{c}$ with analytic coefficients.
We can eliminate $x$ by expanding everything in the Fourier basis. Write
$\nabla_{x}V(x;\mathbf{c})=\sum_{m}c_{m}(2\pi im)e^{2\pi im\cdot x},$
and similarly for $\det H(V)(x;\mathbf{c})$ which becomes a sum $\sum_{m}c_{m}^{(d)}e^{2\pi im\cdot x}$ where $c_{m}^{(d)}$ are homogeneous polynomials of degree $d$ in the coefficients $\{c_{m}\}$ (coming from products of derivatives). The condition that there exists $x\in\mathbb{T}^{d}$ such that all these Fourier series vanish simultaneously is equivalent to the vanishing of all their Fourier coefficients (since the exponentials are linearly independent). More concretely, let $\{e_{k}(x)\}_{k\in\mathcal{K}}$ be a finite basis of trigonometric polynomials of degree $\leq n$ (e.g., $e^{2\pi im\cdot x}$ with $|m|\leq n$). Then $\nabla_{x}V$ and $\det H(V)$ are finite linear combinations of these basis functions. The existence of an $x$ making them zero is equivalent to the existence of $x$ such that a finite system of analytic equations holds. Eliminating $x$ by using resultants (or by noting that the condition that a finite set of analytic functions has a common zero is itself an analytic condition on the coefficients $\mathbf{c}$) yields a finite set of analytic functions $f_{1}(\mathbf{c}),\dots,f_{r}(\mathbf{c})$ such that
$\pi(\mathcal{D})=\{\mathbf{c}\in\mathbb{R}^{N}:f_{1}(\mathbf{c})=\cdots=f_{r}(\mathbf{c})=0\}.$
In other words, $\pi(\mathcal{D})$ is an analytic subset of $\mathbb{R}^{N}$ (in fact, it is a real algebraic set because all equations are polynomial in $\mathbf{c}$ after clearing denominators, but analytic suffices).
We claim that $\pi(\mathcal{D})$ is not the whole space $\mathbb{R}^{N}$. Indeed, take the specific potential $V_{0}(x)=\sum_{j=1}^{d}\cos(2\pi x_{j})$. Its coefficient vector $\mathbf{c}_{0}$ corresponds to a Morse function (the critical points are at $x_{j}=0$ or $1/2$, and the Hessian is non-degenerate). For this $\mathbf{c}_{0}$, there is no $x$ with $\nabla V=0$ and $\det H(V)=0$ simultaneously. Hence $\mathbf{c}_{0}\notin\pi(\mathcal{D})$, so $\pi(\mathcal{D})\neq\mathbb{R}^{N}$.
A proper analytic subset of $\mathbb{R}^{N}$ has Lebesgue measure zero (it is contained in the zero set of a non-zero analytic function, and the zero set of a non-zero analytic function has measure zero). Therefore $\mathcal{B}^{(i)}=\pi(\mathcal{D})$ has measure zero.
(For readers who prefer an even more elementary argument: For each fixed $x$, the conditions $\nabla_{x}V(x;\mathbf{c})=0$ and $\det H(V)(x;\mathbf{c})=0$ define a proper algebraic subset $A_{x}\subset\mathbb{R}^{N}$ (since they are non-trivial linear/polynomial equations). The union $\bigcup_{x\in\mathbb{T}^{d}}A_{x}$ is the set of $\mathbf{c}$ for which there exists some $x$ with degenerate critical point. However, this union is uncountable and could be large; but one can prove that $\bigcup_{x}A_{x}$ is actually a finite union because the equations are analytic and the set of $x$ that can give a solution for some $\mathbf{c}$ is compact and the condition is semi-algebraic. Alternatively, the earlier elimination argument is standard and rigorous.)
@@
Let
$\mathcal{B}^{(ii)}=\{\mathbf{c}\in\mathbb{R}^{N}\mid V(\cdot;\mathbf{c})\text{ lacks unique global extrema}\}.$
We show that $\mathcal{B}^{(ii)}$ has measure zero. By Part 1, for almost every $\mathbf{c}$, $V(\cdot;\mathbf{c})$ is Morse, hence its critical points are isolated. Global extrema are attained at critical points. Uniqueness fails if there exist two distinct critical points $x\neq y$ with $V(x;\mathbf{c})=V(y;\mathbf{c})$.
Define the critical set
@@
For a fixed $\mathbf{c}$ where $V(\cdot;\mathbf{c})$ is Morse, the set $\mathcal{C}_{\mathbf{c}}=\{x:\nabla_{x}V(x;\mathbf{c})=0\}$ is finite and its cardinality is bounded by a constant depending only on the degree $n$. Moreover, the map $\pi:\mathcal{C}\to\mathbb{R}^{N}$ given by $\pi(x,\mathbf{c})=\mathbf{c}$ has the property that for almost every $\mathbf{c}$ the fiber $\pi^{-1}(\mathbf{c})$ is discrete. By the implicit function theorem, $\mathcal{C}$ is a smooth submanifold of $\mathbb{T}^{d}\times\mathbb{R}^{N}$ of dimension NN (the dimension of the parameter space). The exceptional set where this fails is contained in the set of $\mathbf{c}$ where $V(\cdot;\mathbf{c})$ has a degenerate critical point, which has measure zero by Part 1.
Consider the double critical set
@@
which is a smooth manifold of dimension 2Nd2N-d (the constraint $x\neq y$ removes a submanifold of codimension $d$). Define the difference map
@@
The set of parameters for which $V(\cdot;\mathbf{c})$ has non-unique extrema is contained in the projection $\pi(\mathcal{E})$, where $\mathcal{E}=\Xi^{-1}(0)$ and $\pi((x,\mathbf{c}),(y,\mathbf{c}))=\mathbf{c}$.
We claim that $\Xi$ is transverse to $\{0\}$. Indeed, for $((x,\mathbf{c}),(y,\mathbf{c}))\in\mathcal{E}$, a tangent vector $(\dot{x},\dot{\mathbf{c}}_{1},\dot{y},\dot{\mathbf{c}}_{2})$ in $T_{(x,\mathbf{c})}\mathcal{C}\times T_{(y,\mathbf{c})}\mathcal{C}$ satisfies
$D\Xi=\langle\nabla_{x}V(x;\mathbf{c}),\dot{x}\rangle+\langle D_{\mathbf{c}}V(x;\mathbf{c}),\dot{\mathbf{c}}_{1}\rangle-\langle\nabla_{y}V(y;\mathbf{c}),\dot{y}\rangle-\langle D_{\mathbf{c}}V(y;\mathbf{c}),\dot{\mathbf{c}}_{2}\rangle.$
@@
Because $(x,\mathbf{c})$ and $(y,\mathbf{c})$ lie in $\mathcal{C}$, $\nabla_{x}V=\nabla_{y}V=0$, so the terms with $\dot{x},\dot{y}$ vanish. Moreover, since $\mathbf{c}$ is the same for both points, we have $\dot{\mathbf{c}}_{1}=\dot{\mathbf{c}}_{2}=\dot{\mathbf{c}}$ (the tangent space of $\mathcal{C}^{(2)}$ forces this identification). Thus
$D\Xi=\langle D_{\mathbf{c}}V(x;\mathbf{c})-D_{\mathbf{c}}V(y;\mathbf{c}),\dot{\mathbf{c}}\rangle.$
The vectors $D_{\mathbf{c}}V(x;\mathbf{c})=(e^{2\pi im\cdot x})_{|m|\leq n}$ and $D_{\mathbf{c}}V(y;\mathbf{c})=(e^{2\pi im\cdot y})_{|m|\leq n}$ are linearly independent for $x\neq y$ because the exponentials are linearly independent. Hence $D_{\mathbf{c}}V(x;\mathbf{c})-D_{\mathbf{c}}V(y;\mathbf{c})\neq 0$, and by varying $\dot{\mathbf{c}}$ we obtain a surjective derivative onto $\mathbb{R}$. Therefore $\Xi$ is transverse to $\{0\}$, and $\mathcal{E}$ is a smooth submanifold of $\mathcal{C}^{(2)}$ of codimension 11. Consequently,
$\dim\mathcal{E}=\dim\mathcal{C}^{(2)}-1=2N-d-1.$
@@
For sufficiently high degree $n$, we have $N>d+1$, so $\dim\mathcal{E}<N$. The projection $\pi|_{\mathcal{E}}:\mathcal{E}\to\mathbb{R}^{N}$ is smooth. By Sard’s theorem, the set of critical values of $\pi|_{\mathcal{E}}$ has measure zero in $\mathbb{R}^{N}$. The set of parameters $\mathbf{c}$ for which $V(\cdot;\mathbf{c})$ has non-unique extrema is contained in this set of critical values (if $((x,\mathbf{c}),(y,\mathbf{c}))\in\mathcal{E}$ is a regular point of $\pi|_{\mathcal{E}}$, then nearby parameters have distinct values; non-uniqueness can only occur at critical values). Hence $\mathcal{B}^{(ii)}$ has Lebesgue measure zero.